OpenStudy (anonymous):

Indicate which property is illustrated in Step 2. Step 1 6 - 6x - 6 + 7x = (6 - 6x) + (-6 + 7x) Step 2 = (6 - 6x) + (7x - 6) Step 3 = 6 + (-6x + 7x) - 6 Step 4 = 6 + (-6 + 7)x - 6 Step 5 = 6 + x - 6 Step 6 = x + 6 - 6 Step 7 = x + 0 Step 8 = x commutative identity associative distributive

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

would this be commutative because it reflects but at the same time it has parenthesis so would it be associative @jigglypuff314

jigglypuff314 (jigglypuff314):

going from step 1 to step 2 was just flipping two numbers around but it did not use the parenthesis in the switching of the two numbers so still commutative commutative is like: a + b + c = a + c + b while associative is like: (a + b) + c = a + (b + c)

jigglypuff314 (jigglypuff314):

and (a + b) + c = (b + a) + c would be commutative

OpenStudy (anonymous):

ok so i was right just confused about associative @jigglypuff314